# Cost-benefit analysis of an AI-driven operational digital platform for integrated electric mobility, renewable energy, and grid management

### v1

@@ 3.1 ODP system architecture and AI engine
| Perception layer | Collect data on vehicles | EV/ET sensors | BMS sensors, GPS, OCPP-compliant charge monitors |
| Collect data on RES | RES sensors | Pyranometers, anemometers, power meters | |
| Collect data on environmental conditions | Environmental sensor arrays | IoT weather nodes, multi-sensor modules | |
| Collect data on weather conditions | Weather sensors | Integrated weather stations, satellite APIs | |
| Transmit data | Connectivity solutions | 5G, LTE-M/NB-IoT, LoRaWAN, Satellite | |
| Data layer | Receive data | Communication protocols | MQTT, HTTP/HTTPS, OPC-UA, AMQP |
| Communicate data | Message queues | Apache Kafka, RabbitMQ | |
| Store data | Databases | PostgreSQL, MongoDB, InfluxDB, TimescaleDB | |
| Store big data | Data lakes/warehouses | Azure Data Lake, Amazon S3, Hadoop HDFS | |
| Move data | Data integration, ETL | Apache NiFi, Talend, Azure Data Factory, AWS Glue | |
| Middleware layer | Integrate with existing systems | Integration middleware | FIWARE Context Broker, MuleSoft, Apache Camel |
| Provide clean APIs | API management | FIWARE Wilma/Keyrock, Apigee, AWS API Gateway | |
| Integrate with EU data spaces | Data space connectors | FIWARE Connector, IDS Connectors, Gaia-X | |
| Collect market data | Market integration middleware | EPEX SPOT API, Nord Pool API, ENTSO-E API | |
| Business layer (AI Engine Core) | Predict grid load, RES, prices | AI forecasting models | LSTM-RNNs, multi-head attention, NWP features, Azure ML, cloud GPU |
| Optimize EV/ET charging, V2X/V2B | AI optimization, RL agents | MILP/MIQP (CPLEX, Gurobi), DQN/PPO (RL), Python APIs, battery models | |
| Manage grid flexibility | Demand response management | OpenADR 2.0b tools, DR event prediction, load aggregation algorithms | |
| Optimize EV/ET routing | Route optimization engines | Electric VRP solvers, Google OR-Tools, metaheuristics, battery-aware routing | |
| Weather forecast integration | Advanced weather modeling | NWP models (ECMWF, GFS), AI correction (GBM, NN), sensor fusion | |
| Application layer | User interaction | Mobile/web apps | React Native, Flutter, charge planning, fleet dashboards |
| Navigation to charging points | Navigation systems | Google Maps API, Mapbox, HERE, ODP charger data integration | |
| ODP insights | Analytics/visualization | Power BI, Grafana, Tableau, KPI dashboards | |
| Administration tools | Management dashboards | Grafana, Kibana, model drift detection, admin UI |
@@
Note: The ODP architecture is based on the EU-funded BEGONIA project; all layers are described in detail in Deliverable D3.1 [35].
The Perception layer is responsible for real-time data acquisition from a multitude of distributed assets, including EVs, ETs, RES installations (solar PV, wind turbines), charging infrastructure, and environmental sensors. This layer employs various sensors (BMS data, GPS, smart meters, pyranometers, anemometers) and communication technologies (5G, LTE-M, LoRaWAN) to ensure robust and timely data transmission. Data is collected at granular resolutions (e.g., 1-minute intervals for critical parameters). The Data layer manages the ingestion, processing, storage, and secure handling of these voluminous and heterogeneous data streams. It leverages communication protocols (MQTT, OPC-UA), message queuing systems (Apache Kafka), and a hybrid data storage architecture comprising time-series databases (InfluxDB, TimescaleDB) for operational data and data lakes (e.g., Azure Data Lake) for large-scale analytics and AI model training. Data quality validation, anomaly detection, and missing data imputation are integrated processes within this layer to ensure robust inputs for the AI engine.
@@
The Business layer constitutes the AI engine core of the ODP, housing the advanced algorithms and models that drive intelligent decision-making. Key AI components include:
@@
Forecasting models: Long Short-Term Memory (LSTM) based Recurrent Neural Networks (RNNs), enhanced with multi-head attention mechanisms, are employed for forecasting EV/ET charging demand, RES generation, and grid load at various temporal and spatial resolutions (e.g., 15-minute intervals up to 72 hours ahead). These models capture complex temporal dependencies and non-linear relationships inherent in energy data, using inputs like historical load/generation, high-resolution weather forecasts (temperature, irradiance, wind speed, cloud cover from numerical weather prediction models and local sensors), calendar effects (day-of-week, holidays), and real-time market signals. Model training is performed on extensive historical datasets (e.g., 2-3 years of data from ENTSO-E, national TSOs/DSOs, and representative EV/ET usage profiles) using scalable cloud-based GPU resources. Performance is continuously validated through rolling-window cross-validation and backtesting against established baseline methods (e.g., ARIMA, persistence), achieving a target Mean Absolute Percentage Error (MAPE) below 5% for day-ahead RES and load forecasts, demonstrating a significant uplift in predictive power essential for subsequent optimization steps. Energy price forecasting models, potentially hybrid approaches combining statistical models (e.g., ARIMA-GARCH) with machine learning for residual analysis, are also developed to predict day-ahead and intra-day electricity prices.
EV/ET smart charging and V2X dispatch: mixed-integer linear programming (MILP) or mixed-integer quadratic programming (MIQP) formulations optimize charging/discharging schedules for individual vehicles and aggregated fleets. The objective functions are multi-faceted: minimize total energy cost for users, maximize self-consumption of local RES, minimize carbon emissions, reduce peak grid load, and critically, promote battery longevity. Battery degradation is explicitly incorporated into the objective by penalizing charging strategies known to accelerate degradation(e.g., frequent deep discharges, high state-of-charge (SoC) extremes, high C-rates). For instance, additional cost term is included, where represents a battery aging model based on SoC range, C-rates, and temperature, informed by established research [36, 37, 38]. The AI optimizer thus actively minimizes this degradation cost as part of its multi-objective optimization. Constraints include battery SoC limits (), user-defined departure times and required SoC targets, maximum charging/discharging rates, local grid capacity limits, charger availability (from Perception layer), and electricity tariff structures. Commercial solvers like CPLEX or Gurobi are employed via Python APIs. For very large problem instances (thousands of vehicles), decomposition techniques (e.g., column generation) or distributed optimization methods ensure solutions within operational timeframes (minutes for day-ahead scheduling, seconds for real-time adjustments). In addition, V2G/V2B services are managed by reinforcement learning (RL) agents (e.g., Deep Q-Networks or Proximal Policy Optimization). These agents are trained on high-fidelity simulations of real-time energy markets (reflecting bidding rules, settlement periods for frequency regulation services) and grid conditions. The RL reward function incentivizes revenue from ancillary services while safeguarding battery health and user mobility needs. Simulations incorporate historical market data, realistic network constraints, and stochastic elements to approximate real-world uncertainty.
@@
ET route and logistics optimization: Sophisticated vehicle routing problem (VRP) algorithms, specifically tailored for electric fleets (E-VRPTW – Electric VRP with time windows and charging stops), are implemented. These algorithms optimize routes to minimize total operational cost (combining energy consumption, time, and potential battery degradation costs) while considering real-time traffic conditions, terrain (for energy consumption estimation), payload, dynamic weather conditions, charging station availability, charging duration, and potential queuing times at charging points. The integration with battery degradation models ensures that routing decisions do not inadvertently lead to accelerated battery wear. Algorithms are based on advanced metaheuristics (e.g., tabu search, genetic algorithms) or commercial solvers (e.g., Google OR-Tools). The performance is evaluated by metrics such as route efficiency (km/kWh), total travel time, and number of required charging stops, compared to static or less intelligent routing systems.
@@
Demand response (DR) management: AI-driven tools automate the identification and aggregation of flexible EV/ET loads, predict their response potential based on historical patterns and current vehicle states, and manage their participation in DR events or ancillary service markets. This adheres to standards like OpenADR 2.0b. The system uses AI-driven baseline calculation methods to accurately measure DR event effectiveness and optimize dispatch signals to individual vehicles or aggregated charging stations.
@@
These AI models are developed using frameworks like TensorFlow, Keras, and PyTorch, integrating with optimization solvers such as CPLEX or Gurobi via Python APIs. Model training utilizes high-quality historical and real-time data from the Data layer. Crucially, the ODP incorporates continuous operational monitoring, performance validation metrics (e.g., forecast accuracy, optimization solution quality, solver runtime), and scheduled retraining pipelines to proactively address potential data drift, concept drift, or model degradation over time, ensuring sustained AI model efficacy and robust decision-making. Cloud-native scalable infrastructure supports the computational demands of large-scale AI model training and real-time inference.
@@ 3.2 Cost-benefit analysis framework
To evaluate the proposed AI-driven ODP, we employ a structured seven-step CBA framework, guided by European Commission smart grid guidelines [11] and lessons from pilot studies such as Isernia [12]. This framework systematically captures AI-related costs (platform development, sensors and IoT devices, model training and maintenance) and the benefits directly attributable to AI-driven performance (e.g., improved forecast accuracy, optimization savings). The framework starts with a precise definition of the project, its objectives, engineering features, geographic scope, stakeholders, and expected socio-economic impact, and delineation of its functional scope. Next, it maps the enabling technologies (especially AI modules) to the key functions they perform, and links those functions to the spectrum of benefits they generate. Subsequently, each category of benefit is quantified and monetized, and all associated costs (capital and operating) are tallied over the project’s life cycle. Finally, total benefits and costs are compared, and a rigorous sensitivity analysis, including scenario variations and Monte Carlo simulation, is conducted to assess how results vary under key uncertainties (such as AI performance and adoption rates). Table 2 summarizes these steps.
@@
With this framework, Step 1 establishes the project’s scope, objectives, geographic coverage, stakeholders, and counterfactual baseline. Step 2 links each ODP technology layer and AI component to its operational function (data acquisition, integration, forecasting & optimization, decision support). Step 3 identifies how these functions generate economic, reliability, and environmental benefits. In Step 4, each benefit is converted into monetary terms over the project horizon using transparent formulas and country-specific data. Step 5 aggregates all capital and operating expenditures necessary for deployment and operation, including AI model development, sensors, cloud services, and personnel. Step 6 computes net present value (NPV) and benefit–cost ratio (BCR) by discounting all cash flows, reflecting societal time preferences. Finally, Step 7 tests the robustness of the results through scenario variations and Monte Carlo simulation, and identifies key risks with corresponding mitigation strategies. Section 4 applies these steps in detail to the Austria–Hungary–Slovenia case study.
@@ 4 Cost-benefit analysis steps with case study
We now apply the seven-step CBA framework to the AI-driven ODP project, evaluating its deployment across Austria (AT), Hungary (HU), and Slovenia (SI). For the economic appraisal, all future cash flows are expressed in constant 2025 Euros and discounted to present value using a social discount rate of 4%. This choice is consistent with EU recommendations for long-term public infrastructure projects [39]. The results highlight the value created by the ODP’s AI-driven features and provide insight into the project’s economic viability under various scenarios.
@@
The project under consideration is the deployment and operation of the AI-driven ODP for the integrated management of EVs, ETs, RES, and the electric grid across three countries: Austria, Hungary, and Slovenia. The primary objective of this project is to leverage real-time data and advanced AI (for forecasting and control) to enhance grid flexibility, reduce energy procurement and operational costs, minimize greenhouse gas emissions and local air pollutants, improve power system reliability, and enable seamless interoperability between transport electrification and power system operations. The implementation scale is national/regional within these countries, which have been chosen for their diverse EV/ET adoption rates, renewable energy capacities, and grid characteristics, thereby providing a representative range of EU contexts for analysis.
@@
Key engineering features of the ODP critical to the project include: (a) AI-powered predictive algorithms for demand forecasting (EV/ET charging, overall grid load), RES generation forecasting, and energy price forecasting. These provide high accuracy (e.g., day-ahead load and RES forecasts with MAPE ), superior to conventional methods, directly translating to improved operational decisions and cost savings. (b) Optimization engines (MILP-based solvers and RL agents) for smart charging scheduling, V2X services, electric fleet logistics (routing and charging), and demand response dispatch. These engines optimize multi-objective functions, including cost reduction, emissions reduction, and battery health preservation, thus enabling coordinated control strategies that balance the grid and maximize economic returns. (c) IoT infrastructure comprising a dense network of sensors and smart meters to provide granular, real-time monitoring of vehicles(battery SoC, location, charging status), charging stations (availability, power flow), grid nodes (voltage, load), and RES installations (output, status). This high-resolution data collection is foundational for the AI engine’s situational awareness. (d) High-speed, reliable communication networks (5G, LTE-M) ensuring low-latency data transfer and control signal dissemination. This is critical for real-time AI inference and control, allowing the ODP to rapidly respond to changing conditions (such as sudden RES output fluctuations or grid contingencies). (e) Robust middleware integration with existing energy management systems (utility SCADA/DMS/EMS) and market platforms. By using standardized interfaces and data exchange protocols, the ODP can seamlessly interoperate with legacy systems and participate in energy markets (for energy trading and ancillary services), thus embedding the platform within the broader energy ecosystem.
The primary stakeholders expected to benefit from this ODP include: EV owners and ET fleet operators (through optimized charging costs, reduced range anxiety, improved charging availability, extended battery lifespan, and efficient routing); charging point operators (CPOs) (via enhanced asset utilization and new service offerings); distribution and transmission system operators (DSOs/TSOs) (through reduced grid congestion, deferred network reinforcement investments, improved grid stability, and enhanced situational awareness due to AI-driven insights); renewable energy producers (via reduced curtailment and better market integration of variable generation); and society at large (through reduced carbon emissions, improved air quality, and enhanced energy security).
The baseline scenario (counterfactual case without the ODP) assumes that EV and ET adoption and renewable deployment continue to grow in the three countries, but in a conventional manner: EV charging remains largely uncoordinated (users charge at will, leading to sharper peaks and low overnight utilization), V2G and other vehicle-to-grid services have minimal participation, RES curtailment issues persist at times of excess generation, and grid operators have limited visibility/control over distributed EV chargers and other flexible resources. Essentially, the baseline reflects current practices where the transport and energy sectors operate with only rudimentary interaction and rely on existing, less sophisticated forecasting and control tools. The cost-benefit analysis will evaluate the incremental impacts of implementing the ODP relative to this baseline, thereby isolating the net benefits attributable to the platform’s advanced AI-driven coordination.
@@
In this step, each core technology component of the ODP, particularly its AI modules as detailed in Section 3.1, is systematically mapped to the specific operational functions it enables. The ODP integrates a comprehensive suite of technologies (sensors, communication networks, data platforms, AI algorithms, user interfaces), which together deliver several key functions for the system’s operation:
(A) Advanced data and information collection & processing: Real-time gathering, validation, and processing of granular data from EVs/ETs (e.g., battery SoC, location, charging patterns), RES installations (generation output, availability), grid sensors (voltage levels, loading), charging infrastructure status, and external sources (weather, market market prices). This function is enabled by IoT sensors, connectivity solutions, and the ODP’s Data layer; it ensures that the AI engine has accurate and timely information about the system’s state at all times.
(B) Enhanced interoperability and system integration: Seamless data exchange and coordinated operation among diverse systems and stakeholders (vehicles, grid operators, CPOs, energy markets, end-users). This is facilitated by the Middleware layer and use of standardized communication protocols and interfaces (e.g., OCPP for charging, OpenADR for demand response signals, API gateways for market data). Enhanced interoperability ensures that decisions made by the AI core can be effectively communicated and executed across different subsystems in real time.
@@
i) Intelligent forecasting: The AI models (LSTMs, ML models, etc., as described in Section 3.1) provide highly accurate predictions for EV/ET charging demand, RES generation, grid load, and electricity prices. These forecasts improve upon traditional methods, reducing uncertainty and prediction error in the system’s operation.
ii) Smart load management and grid balancing: AI optimization algorithms (MILP, RL, E-VRPTW, as detailed in Section 3.1) enable dynamic smart charging/discharging of EVs/ETs, optimal dispatch of V2X services for grid support, dynamic EV/ET routing considering energy and battery health, and coordinated demand response actions. These control decisions continuously balance supply and demand, flatten load curves, alleviate grid constraints, and maximize the use of renewable energy.
(D) Decision support and automation: User-friendly interfaces (mobile apps, dashboards) provide actionable insights and recommendations to end-users (e.g., optimal charging times) and system operators (e.g., congestion alerts and recommended mitigation actions). Automated control systems execute optimized schedules and dispatch signals directly generated by the AI engine, ensuring real-time response and reducing manual intervention.
Table 3 provides a high-level mapping of the ODP’s major technological assets and AI components to these functional categories A–D. For instance, IoT sensors and connectivity primarily support data collection (Function A). The integration middleware and communication protocols enable Function B. The AI forecasting models and optimization algorithms directly drive Function C (intelligent forecasting and control), while the user interfaces and automated control modules deliver Function D (decision support and action). This detailed mapping ensures that all critical functions necessary for the ODP’s success are supported by specific, well-defined technological capabilities, with the core AI engine acting as the central intelligence orchestrating these functions.
@@
i) Economic benefits: Direct financial gains or cost savings, encompassing reduced operational expenditures for grid operators and end-users, new revenue streams from energy services (e.g., ancillary services, energy arbitrage), enhanced energy efficiency, and optimized asset utilization.

### v2

@@ 3.1 ODP system architecture and AI engine
[…]
@@
Note: The ODP architecture is based on the EU-funded BEGONIA project; all layers are described in detail in Deliverable D3.1 Mombartz-Brændgaard et al. [2024].
Note that the scope here is to assess the economic and financial viability of the proposed AI-driven ODP via a transparent CBA framework. This means we do not empirically validate or optimize the AI modules or the core architecture presented in Table LABEL:tab_architecture. Accordingly, any performance figures (e.g., forecasting accuracy or coordination gains) are treated based on the literature values of the platform when deployed in the implementation phase, rather than as case-study results. The requirements and benefits of each architectural layer should be read together with the technological diagram in Figure 1 to understand the platform-level interconnectedness. The CBA then evaluates the economic and financial benefits and expenditures required to implement this ODP using the case study. Below are the details of each layer:
The Perception layer is responsible for real-time data acquisition from a multitude of distributed assets, including EVs, ETs, RES installations (solar PV, wind turbines), charging infrastructure, and environmental sensors. This layer employs various sensors (BMS data, GPS, smart meters, pyranometers and anemometers) and communication technologies (5G, LTE-M, LoRaWAN) to ensure robust and timely data transmission. Data is collected at granular resolutions (e.g., 1-minute intervals for critical parameters).
The Data layer manages the ingestion, processing, storage, and secure handling of these voluminous and heterogeneous data streams. It leverages communication protocols (MQTT, OPC-UA), message queuing systems (Apache Kafka), and a hybrid data storage architecture comprising time-series databases (InfluxDB, TimescaleDB) for operational data and data lakes (e.g., Azure Data Lake) for large-scale analytics and AI model training. Data quality validation, anomaly detection, and missing data imputation are integrated processes within this layer to ensure robust inputs for the AI engine.
@@
The Business layer constitutes the core of the AI engine in the ODP, housing advanced algorithms and models that drive intelligent decision-making. Key AI components include:
@@
Forecasting models: Long Short-Term Memory (LSTM) based Recurrent Neural Networks (RNNs), enhanced with multi-head attention mechanisms, are employed for forecasting EV/ET charging demand, RES generation, and grid load at various temporal and spatial resolutions (e.g., 15-minute intervals up to 72 hours ahead). These models capture complex temporal dependencies and non-linear relationships inherent in energy data, using inputs, such as historical load/generation, high-resolution weather forecasts (temperature, irradiance, wind speed, cloud cover from numerical weather prediction models, and local sensors), calendar effects, and real-time market signals. Model training is performed on extensive historical datasets, demonstrating a significant uplift in predictive power essential for subsequent optimization steps. Energy price forecasting models with machine learning for residual analysis, are alsoincluded in the core ODP architecture to predict day-ahead and intra-day electricity prices.
EV/ET smart charging and V2X dispatch: for this service, mixed-integer linear programming (MILP) or mixed-integer quadratic programming (MIQP) formulations optimize charging/discharging schedules for individual vehicles, and aggregated fleets can be augmented. Within the ODP, smart charging and V2X coordination are treated as benefit channels. The platform schedules charging and limited vehicle-to-grid export to align with RES availability, tariff signals, and grid limits, while respecting operational requirements (e.g., departure times and minimum state-of-charge). Any battery-health safeguards (e.g., avoiding extreme SoC or excessive C-rates) are handled as practical constraints rather than as a separate modeling contribution. The CBA captures the incremental value of this function through measurable KPI deltas, such as reduced peak demand, avoided curtailment, lower imbalance costs, and modest maintenance savings, relative to a clearly defined without-ODP baseline. In relation to the ODP AI-engine, the V2G/V2B services are managed by reinforcement learning (RL) agents (e.g., Deep Q-Networks or Proximal Policy Optimization). These agents are trained on high-fidelity simulations of real-time energy markets (reflecting bidding rules, settlement periods for frequency regulation services) and grid conditions. The RL reward function incentivizes revenue from ancillary services while safeguarding battery health and user mobility needs. Simulations incorporate historical market data, realistic network constraints, and stochastic elements to approximate real-world uncertainty.
@@
ET route and logistics optimization: Sophisticated vehicle routing problem (VRP) algorithms, specifically tailored for electric fleets (E-VRPTW – Electric VRP with time windows and charging stops), are implemented. These algorithms optimize routes to minimize total operational cost (combining energy consumption, time, and potential battery degradation costs) while considering real-time traffic conditions, terrain (for energy consumption estimation), payload, dynamic weather conditions, charging station availability, charging duration, and potential queuing times at charging points. The integration with battery degradation models ensures that routing decisions do not inadvertently lead to accelerated battery wear.
@@
Demand response (DR) management: AI-driven tools automate the identification and aggregation of flexible EV/ET loads, predict their response potential based on historical patterns and current vehicle states, and manage their participation in DR events or ancillary service markets. This adheres to standards like OpenADR 2.0b. The system utilizes AI-driven baseline calculation methods to accurately measure the effectiveness of DR events and optimize dispatch signals for individual vehicles or aggregated charging stations.
@@
These AI models are developed using frameworks such as TensorFlow, Keras, and PyTorch, integrating with optimization solvers, such as CPLEX or Gurobi via Python APIs. Model training utilizes high-quality historical and real-time data from the Data layer. Crucially, the ODP incorporates continuous operational monitoring, performance validation metrics (e.g., forecast accuracy, optimization solution quality, solver runtime), and scheduled retraining pipelines to proactively address potential data drift, concept drift, or model degradation over time, ensuring sustained AI model efficacy and robust decision-making. Cloud-native scalable infrastructure supports the computational demands of large-scale AI model training and real-time inference.
@@
Figure 1 illustrates the interconnection of the five principal layers of the ODP. The perception layer collects real-time data from EV, ET, and RES sensors, as well as environmental and weather monitoring devices, which are transmitted via communication protocols and connectivity solutions. The middleware layer manages this data through message queues, ETL pipelines, and integration middleware, enabling secure exchange with external data spaces and market platforms. Processed data are then stored in data lakes and databases, forming the foundation for the service layer, where AI models for grid load, weather forecasting, demand response, and route optimization operate. The outputs feed into user-facing services such as mobile applications, dashboards, and analytics tools, which constitute the business layer by supporting operational decisions, market integration, and stakeholder engagement. This layered and interconnected design ensures scalability, interoperability, and coordinated optimization across EVs, ETs, RES, and the grid.
While the ODP–CBA relies on country-specific inputs (RES generation/forecasts, EV/ET fleets and charging schedules, grid constraints and states, cross-border flows), these data are increasingly standard in Europe due to smart-grid deployments, national data hubs, and harmonized interfaces. The framework is designed to plug into these interoperable data rails, making large-scale deployment practical where electrification and data-sharing initiatives are underway.
@@ 3.2 Cost-benefit analysis framework
To evaluate the proposed AI-driven ODP, we employ a structured seven-step CBA framework, inspired by the European Commission smart grid guidelines [EU Commission JRC, 2012] and lessons from pilot studies such as Isernia [Stromsather et al., 2018]. The proposed framework systematically captures AI-related costs (platform development, sensors and IoT devices, model training and maintenance) and the benefits directly attributable to AI-driven performance (e.g., improved forecast accuracy, and optimization savings). The framework begins with a precise definition of the project, its objectives, engineering features, geographic scope, stakeholders, and expected socio-economic impact, and delineation of its functional scope. Next, it maps the enabling technologies (especially AI modules) to the key functions they perform, and links those functions to the spectrum of benefits they generate. Subsequently, each category of benefit is quantified and monetized, and all associated costs (capital and operating) are tallied over the project’s life cycle. Finally, total benefits and costs are compared, and a rigorous sensitivity analysis, including scenario variations and Monte Carlo simulation, is conducted to assess how results vary under key uncertainties. Table 2 summarizes these steps.
@@
With this framework, Step 1 establishes the project’s scope, objectives, geographic coverage, stakeholders, and counterfactual baseline. Step 2 links each ODP technology layer and AI component to its operational function (data acquisition, integration, forecasting & optimization, and decision support). Step 3 identifies how these functions generate economic, reliability, and environmental benefits. In Step 4, each benefit is converted into monetary terms over the project horizon using transparent formulas and country-specific data. Step 5 aggregates all capital and operating expenditures necessary for deployment and operation, including AI model development, sensors, cloud services, and personnel. Step 6 computes net present value (NPV) and benefit–cost ratio (BCR) by discounting all cash flows, reflecting societal time preferences. Finally, Step 7 tests the robustness of the results through scenario variations and Monte Carlo simulation, and identifies key risks and corresponding mitigation strategies.
@@ 4 Cost-benefit analysis steps with case study
We now apply the seven-step CBA framework to the AI-driven ODP project, evaluating its deployment across Austria (AT), Hungary (HU), and Slovenia (SI). For the case study, i) the choice of the three countries is based on stakeholder analysis, the cross-sectoral and cross-border nature of the digital platforms. ii) The chosen countries represent different RES projection capacities, EV/ET penetrations, and input parameter prices. For data and parameter values, three assumptions are considered: iii) country-specific assumptions, technology-specific assumptions, and economic assumptions. All input assumptions, input parameters, and sources are provided in Tables (5, 10, 10, and 13). For the economic appraisal, all future cash flows are expressed in constant 2025 Euros and discounted to present value using an economic discount rate of 4%. [European Commission Directorate-General for Regional and Urban Policy, 2021]. The project’s primary objective is to use real-time data and advanced AI models to optimize system operations. In doing so, it will enhance grid flexibility, lower energy procurement and operating costs, cut greenhouse-gas and local-air-pollutant emissions, improve power-system reliability, and enable seamless interoperability between transport electrification and grid operations.
@@
The project under consideration is the deployment and operation of the AI-driven ODP for the integrated management of EVs, ETs, RES, and the electric grid across three countries. The ODP’s primary objective is to use real-time data and advanced AI models to optimize system operations. In doing so, it will enhance grid flexibility, lower energy procurement and operating costs, cut greenhouse-gas and local-air-pollutant emissions, improve power-system reliability, and enable seamless interoperability between transport electrification and grid operations. The implementation scale is national/regional within these countries, which have been chosen for their diverse EV/ET adoption rates, renewable energy capacities, and grid characteristics, thereby providing a representative range of EU contexts for analysis.
@@
Key engineering features of the ODP critical to the project include: i) AI-powered predictive algorithms for demand forecasting (EV/ET charging, overall grid load), RES generation forecasting, and energy price forecasting. ii) Optimization engines for smart charging scheduling, V2X services, electric fleet logistics (routing and charging), and demand response dispatch. These engines optimize multi-objective functions, including cost reduction, emissions reduction, and battery health preservation, thus enabling coordinated control strategies that balance the grid and maximize economic returns. iii) IoT infrastructure comprising a dense network of sensors and smart meters to provide granular, real-time monitoring of vehicles, charging stations, grid nodes, and RES installations. This high-resolution data collection is foundational for the AI engine’s situational awareness. iv) High-speed, reliable communication networks ensuring low-latency data transfer and control signal dissemination. This is critical for real-time AI inference and control, allowing the ODP to rapidly respond to changing conditions. v) Robust middleware integration with existing energy management systems (utility SCADA/DMS/EMS) and market platforms. By using standardized interfaces and data exchange protocols, the ODP can seamlessly interoperate with legacy systems and participate in energy markets, thus embedding the platform within the broader energy ecosystem.
As a stakeholder engagement, we conduct structured interviews and workshops with TSOs/DSOs, CPOs, fleet operators, and regulatory authorities to elicit requirements and validate governance, relevance, and technical aspects of the ODP. These assessments informed functional and data-sharing needs, adoption and usage assumptions, and implementation risks spanning governance, interoperability, and regulatory readiness.
Figure 2 summarizes a stakeholder-informed assessment across regulatory, technical, governance, economic/financial, and relevance dimensions for the three case-study countries, providing contextual evidence on organizational readiness and adoption feasibility that complements the quantitative CBA.
The benchmark for the CBA framework is the without-ODP system representing the existing system. The current system is largely uncoordinated in terms of EV/ET charging, has minimal V2G capabilities, and involves residual RES curtailment during surplus periods, as well as limited operator visibility/control due to legacy forecasting and tools. All reported benefits are then incremental relative to this benchmark. For benefit stream at time ,
where denotes the physical metric that is monetized and is annual benefits aggregate over and across countries. Assumptions regarding adoption, charging patterns, RES integration, and market conditions were informed by surveys and interviews with TSOs, DSOs, CPOs, fleet operators, and regulators in Austria, Hungary, and Slovenia.
Participation by EV drivers and fleet operators in smart charging and demand response is modeled explicitly as a scenario input. The CBA assumes rational, benefit-maximizing behavior within country-specific policies and grid codes (EV/ET, RES, network operation). Non-technical barriers (heterogeneous charging habits, DR fatigue, data-sharing and regulatory frictions) are not modeled as separate processes; instead, they are reflected through conservative/central/optimistic participation ranges and assessed qualitatively in the stakeholder-informed multi-criteria evaluation (regulatory, governance, relevance; Figure 2).
@@
In this step, each core technology component of the ODP, particularly its AI modules as detailed in Section 3.1, is systematically mapped to the specific operational functions it enables. The ODP integrates a comprehensive suite of technologies, which together deliver several key functions for the system’s operation:
(A) Advanced data and information collection & processing: Real-time gathering, validation, and processing of granular data from EVs/ETs, RES installations, grid sensors, charging infrastructure status, and external sources (weather, market prices). This function is enabled by IoT sensors, connectivity solutions, and the ODP’s Data layer, ensuring that the AI engine has accurate and timely information about the system’s state at all times.
(B) Enhanced interoperability and system integration: Seamless data exchange and coordinated operation among diverse systems and stakeholders (vehicles, grid operators, CPOs, energy markets, end-users). This is facilitated by the Middleware layer and use of standardized communication protocols and interfaces. Enhanced interoperability ensures that decisions made by the AI core can be effectively communicated and executed across different subsystems in real time.
@@
i) Intelligent forecasting: The AI models (LSTMs, ML models, etc.,) provide highly accurate predictions for EV/ET charging demand, RES generation, grid load, and electricity prices. These forecasts improve upon traditional methods, reducing uncertainty and prediction error in the system’s operation.
ii) Smart load management and grid balancing: AI optimization algorithms (MILP, RL, E-VRPTW) enable dynamic smart charging/discharging of EVs/ETs, optimal dispatch of V2X services for grid support, dynamic EV/ET routing considering energy and battery health, and coordinated demand response actions. These control decisions continuously balance supply and demand, flatten load curves, alleviate grid constraints, and maximize the use of renewable energy.
(D) Decision support and automation: User-friendly interfaces (mobile apps, dashboards) provide actionable insights and recommendations to end-users (e.g., optimal charging times) and system operators (e.g., congestion alerts and recommended mitigation actions). Automated control systems execute optimized schedules and dispatch signals directly generated by the AI engine, ensuring real-time responses and reducing manual intervention.
Table 3 provides a high-level mapping of the ODP’s major technological assets and AI components to these functional categories A–D. For instance, IoT sensors and connectivity primarily support data collection (Function A). The integration middleware and communication protocols enable Function B. The AI forecasting models and optimization algorithms directly drive Function C (intelligent forecasting and control), while the user interfaces and automated control modules deliver Function D (decision support and action). This detailed mapping ensures that all critical functions necessary for the ODP’s success are supported by specific, well-defined technological capabilities, with the core AI engine acting as the central intelligence that orchestrates these functions.
@@
i) Economic benefits: Direct financial gains or cost savings, encompassing reduced operational expenditures for grid operators and end-users, new revenue streams from energy services, enhanced energy efficiency, and optimized asset utilization.